OpenStudy (anonymous):

Determine the perimeter of the triangle whose sides have the midpoints (1,0), (2, -1) and (3,1)

OpenStudy (anonymous):

where did the 2 came from? i mean the constant.

OpenStudy (zzr0ck3r):

(1,0), (2, -1) and (3,1) first two sqrt((-1-0)^2+(2-1)^2) = sqrt(2) second 2 sqrt((1-1)^2+(3-2)^2) = 1 last and first one sqrt((1-0)^2+(3-1)^2) = sqrt(5) so sqrt(5) +sqrt(2) + 1

OpenStudy (anonymous):

owww. okok. thanks :))

OpenStudy (anonymous):

can you also help me find the vertices of those midpoints.

OpenStudy (anonymous):

@zzr0ck3r can you also help me find the vertices of those midpoints.

OpenStudy (kropot72):

A line segment that joins the mid-points of the sides of a triangle is parallel to the third side. Therefore you can use two of the given mid-points to find the slope of the segment that joins the mid points using the formula \[m=\frac{y _{1}-y _{2}}{x _{1}-x _{2}}\] then using this slope and the third mid-point you can find the equation of the line that contains a side of the triangle using the point-slope form of the equation of a line: \[y-y _{1}=m(x-x _{1})\]

ganeshie8 (ganeshie8):

yea, a thing to note for the initial question of finding perimeter : What zz solved was the perimeter of triangle formed by joining midpoints. You need to take twice of that for the actual perimeter of the triangle

OpenStudy (anonymous):

so if the perimeter of the midpoints is n should i multiply it by 2??

ganeshie8 (ganeshie8):

Yes ! the side joining midpoints is half the parallel side

OpenStudy (kropot72):

Then you need to do the above process once more using a different set of mid-points. Now you have have two equations forming a system of equations with a solution set that is the ordered pair that defines one vertex of the triangle.

OpenStudy (anonymous):

ok.thanks for the help :))

ganeshie8 (ganeshie8):

:) continue wid finding vertices... it seems complex

OpenStudy (anonymous):

Yes, multiply by 2. It's an interesting effect when you take a triangle and draw an triangle inside it at the outside triangle's midpoints: You effectively make a similar triangle as the original, except that it's rotated 180 degrees and is half the size. 

OpenStudy (anonymous):

And oh, I forgot, when you draw that inner triangle, all 4 resultant triangles are congruent.